# Volatile-rich Sub-Neptunes as Hydrothermal Worlds: The Case of K2-18 b

### v1

@@ Appendix C Detailed Constraints on Hydrogen Molality
Here, we provide a detailed explanation of how the hydrogen molality ($m_{\text{H}_{2}}$) range used in our modeling was derived. To determine a reasonable range of $m_{\text{H}_{2}}$ values for our modeling, we first turn toward planet interior structure models to provide a range of possible compositions that match the mass and radius of K2-18 b. This approach provides an independent constraint on the masses of H-He and water, which can be used to estimate the equivalent molality of H2 for the bulk planet. We assume that this value applies to a supercritical water ocean.
[…]
@@ Appendix D Dilute Solution Assumptions in the DEW model
Our treatment of water as a dilute solution involves several assumptions that are implicit in our choice of standard states. A standard state is a set of conditions for a species with a fixed chemical composition, serving as a reference state for calculating thermodynamic properties. Often, the standard state is a hypothetical state that may not be physically possible but remains useful in extrapolating thermodynamic properties. We can relate the activities, molalities, and mole fractions through the definitions of standard states from the following equations, adopted from Sverjensky (2019).
@@
where ${\gamma}_{j}$ and ${m}_{j}$ refer to the activity coefficient and molality of species $j$ and ${\gamma}^{0}_{j}$ and ${m}^{0}_{j}$ refer to the standard state values (Anderson, 2005). We adopt a Henryan standard state (i.e., following Henry’s Law) known as the “ideal one molal state” for aqueous species. This is a commonly used hypothetical standard state that describes an ideal solution at infinite dilution with ${m}^{0}_{j}$ = 1.0. By definition, real aqueous species cannot physically exist in this state. Under this condition, since ${m}^{0}_{j}$ = 1.0, ${\gamma}^{0}_{j}$ = 1.0 (see Figure 5 of Sverjensky (2019)), As a result, Equation D1 becomes,
@@
The species of interest in this paper (CH4, CO2, and CO) are neutral, nonpolar molecules, and while their activity coefficients may depart from unity, the ratio of their activity coefficients tends to cancel out. As a result, the ratios of their activities can be approximated as,
$\frac{{a}_{j}}{{a}_{i}}=\frac{{m}_{j}}{{m}_{i}}$ (D4)
@@
where ${\lambda}_{\text{H}_{2}\text{O}}$ is a rational activity coefficient. In the present system, water behaves ideally and ${\lambda}_{\text{H}_{2}\text{O}}$rightarrow 1.0 as ${X}_{\text{H}_{2}\text{O}}$rightarrow 1.0, and therefore,
@@
The results of Equations D6 and D4 are hence substituted into Equations 2 and 4.
@@ Appendix E Generating the Pressure-Temperature Profiles
Care was taken in proximity to the liquid-vapor phase boundary, where a discontinuous jump in the adiabatic temperature gradient can produce errors in the data interpolation. Integration was stopped when the temperature was within 10 K of the phase boundary, and the profiles were then linearly interpolated (in logarithmic pressure space) towards the phase boundary. Warmer adiabats that do not intersect this boundary will have hydrogen in much lower concentrations in the upper atmosphere (at maximum 2% by mass, see Section 2.2), since in this case there is no cold trapping of water and the hydrogen will be fully miscible with supercritical water (Soubiran & Militzer, 2015), which is not consistent with the non-detection of water in K2-18 b’s atmosphere (Madhusudhan et al., 2023b).
@@ Appendix F Parameters for Equilibrium Calculations
The equilibrium constants and Henry’s Law constants in Equations 2, 4, 6, 8, and 10 are obtained with the DEW model, which uses a series of Helgeson-Kirkham-Flowers (HKF) equations of state for aqueous species (Shock & Helgeson, 1988, 1990; Sverjensky et al., 2014) and an equation of state by Zhang & Duan (2005) for water. For convenience, the constants are then fitted with the logarithmic equation and associated values in Table 1. The equilibrium constants in Equations 2 and 4 are fitted on a linear grid of temperatures between 647.15 K - 1273.15 K at every 1 K following the lower limit, resembling the ocean temperature, at each surface pressure (1000 bar, 5000 bar, 10,000 bar).
The Henry’s Law constants are initially fitted on a linear grid of temperatures between 444.15 K to 623.15 K at every 1 K. The lower limit represents the lowest temperature at which our P-T profiles cross the saturation boundary, while the upper limit represents DEW’s calculation limit at 350C (623.15 K). To account for our P-T profiles with saturated temperatures reaching up to 647.15 K, we extrapolate the Henry’s Law constants using the logarithmic equation shown in Table 1 based on the fitting between 444.15 K - 623.15 K. Here, the corresponding pressures are the water saturated vapor pressures, referred to as “Psat Conditions”. Although this extrapolation introduces some uncertainty, the fitting achieves R2 values of at least 0.996, demonstrating a good fit to the data. Figure F.1 shows the trend of Henry’s Law constants across the range of saturated temperatures.
Table 1: Summary of Key Chemical Reactions and Parameters to Calculate Equilibrium Constants
@@ A B C
Henry’s Law for Methane CH4(g) $\rightarrow$ CH4(aq) Psat Conditions 0.997 -120.131 7106.969 38.224
Henry’s Law for Carbon Dioxide CO2(g) $\rightarrow$ CO2(aq) Psat Conditions 0.996 -76.414 4775.558 24.016
Henry’s Law for Carbon Monoxide CO(g) $\rightarrow$ CO(aq) Psat Conditions 0.998 -121.555 6977.213 38.883
Note. — This table presents the parameters for various chemical reactions. The constants A, B, and C are used in the logarithmic equation $\log(K)=A+\frac{B}{T}+C\log(T)$, where T is the absolute temperature.
Figure F.1: Linearly fitted Henry’s Law constants for CH4, CO2, CO across saturated temperatures between 444.15 K to 623.15 K represented by solid lines, with extrapolated values between 624.15 K to 647.15 K represented by dashed lines.
@@ Appendix G Caveats
To address the challenges of modeling K2-18 b’s interior, we employed a geochemical model to investigate the chemistry of a hypothetical supercritical water ocean on K2-18 b. However, the existing DEW model used here is limited to water-rich fluids (see the assumptions in Appendix D) and cannot fully account for scenarios with higher hydrogen molalities. It is possible that K2-18 b could possess significantly higher hydrogen molalities than those modeled here. An upper limit for $m_{\text{H}_{2}}$ can be inferred from JWST-derived atmospheric abundances reported by Madhusudhan et al. (2023b). The volume mixing ratios of CH4 and CO2, each at approximately 1%, provide a lower bound on the bulk volatile content, thus constraining the upper limit on the hydrogen content. Using water as a proxy for all volatiles, this corresponds to a H2 molality upper limit of 1300 $\cdot$(kg H2O)-1. Such a high H2 molality would represent an extremely water-poor fluid, making it incompatible with the DEW model. To investigate these possible H2-rich scenarios, photochemical models (Yang & Hu, 2024) and/or molecular dynamics (MD) simulations such as those used in Gupta et al. (2024) may be more appropriate.
@@
This caveat also affects our P-T profiles and the saturation boundary in Figure 2, and consequently their intersections, where we performed Henry’s law calculations. In this work, both the P-T profiles and the saturation boundary were constructed for pure water, or $X_{\text{H}_{2}\text{O}}$ = 100%. However, K2-18 b’s lower atmosphere has $X_{\text{H}_{2}\text{O}}<100$. To estimate how different H2-H2O mixtures might influence the saturation boundary – and, by extension, the crossing temperatures and pressures – we examined several $X_{\text{H}_{2}\text{O}}$ cases. For $X_{\text{H}_{2}\text{O}}$ = 99% and $X_{\text{H}_{2}\text{O}}$ = 90%, the saturation boundaries shifted by less than 10 K. For $X_{\text{H}_{2}\text{O}}$ = 50%, the maximum shift was approximately 50 K. These temperature changes are not significant enough to greatly impact the results of Henry’s law calculations. However, future work should incorporate P-T profiles and saturation boundaries that account for the effects of species other than water to achieve a more self-consistent model.
@@
We assume that gaseous calculations that are performed at the interface can be directly compared to the JWST observations. However, one notable exception may be CO2, which can remain in a kinetically controlled balance with quenched CO and H2O at altitudes above the CO-CH4-H2O quench point. In that situation, the CO2 mixing ratio can often increase with altitude above the quench point, resulting in a CO2 mole fraction at higher altitudes in excess of expectations based on its deeper equilibrium abundance. Further disequilibrium chemistry gas-phase modeling may be needed to test this scenario.

### v2

@@ Appendix C Detailed Constraints on Hydrogen Molality
Here, we explain how the hydrogen molality ($m_{\text{H}_{2}}$) range used in our modeling was derived. To determine a reasonable range of $m_{\text{H}_{2}}$ values for our modeling, we first turn toward planet interior structure models to provide a range of possible compositions that match the mass and radius of K2-18 b. This approach provides an independent constraint on the masses of H-He and water, which can be used to estimate the equivalent molality of H2 for the bulk planet. We assume that this value applies to a supercritical water ocean.
The radius of K2-18 b can be reproduced by end-member cases where the interior is either i) an Earth-like core with a pure H-He envelope and no water, which corresponds to a hydrogen molality ($m_{\text{H}_{2}}$) of infinity or ii) a nearly 100% water body with no core and negligible amounts of H-He, which leads to $m_{\text{H}_{2}}\approx 0$ (Madhusudhan et al., 2020).
A lower limit on the H2 molality can be inferred from the ice-to-rock ratio accreted by the planet. If the composition of the interior is inherited from the protoplanetary disk (PPD), the latter end member described above is a highly unlikely scenario. The ice-to-rock ratio in a planet would be controlled by the composition of pebbles/planetesimals in the disk, which is typically made of a mixture of 50% ice and 50% rock (iron and silicates) by mass (Drazkowska & Alibert, 2017; Bitsch & Battistini, 2020; Schneider & Bitsch, 2021; Aguichine et al., 2022; Aguichine & Mousis, 2024). Therefore, we place an upper limit on the ice-to-rock content in K2-18 b as 1:1 by mass, similar to the inferred compositions of the icy moons of the Solar System (Sotin et al., 2007). With this upper limit on the bulk water content, K2-18 b requires a minimal amount ($0.02\%$) of H-He to match its radius (see right panel of Figure 3 from Madhusudhan et al. (2020)). Such a composition would correspond to a lower limit $m_{\text{H}_{2}}=0.14$ $\cdot$(kg H2O)-1.
Furthermore, sub-Neptunes should have a limit to how much hydrogen they can accrete during formation. The study of Ginzburg et al. (2016) investigated the maximum hydrogen mass fraction retained by the planet according to the core accretion formalism and after an early stage of boil-off. Using their Equation 24, we estimate that the maximum H-He mass fraction accreted by K2-18 b is $\sim 2\%$. However, applying this limit to the interior modeling from Madhusudhan et al. (2020) may or may not constitute an additional constraint depending on the assumptions of the interior model, namely the intrinsic temperature of the planet $T_{\mathrm{int}}$. For $T_{\mathrm{int}}=25$ K, an upper limit of 2% on the H-He will lead to a bulk composition of 2% H-He and 18% H2O. However, for $T_{\mathrm{int}}=50$ K, only 1.5% of H-He is required to match the bulk density of K2-18 b, meaning that a pure H-He atmosphere cannot be excluded with arguments from interior and formation modeling. Nevertheless, pure H-He envelopes are considered unrealistic, as the hydrogen is expected to react with the FeO component of the silicate mantle to produce H2O (Kite et al., 2020; Schlichting & Young, 2022) and early accretion of planetesimals contribute to the atmosphere’s inventory of metals (Fortney et al., 2013; Mol Lous et al., 2024). A bulk composition of 2% H-He and 18% H2O leads to $m_{\text{H}_{2}}=42$ $\cdot$(kg H2O)-1. In addition, Yang & Hu (2024) derived a bulk envelope H2O/H2 ratio through gas-phase chemistry, finding that a scenario in which K2-18 b has 10×$\times Z_{\odot}$ with H2O/H2 = 50/50 is most consistent with observations. This ratio translates to a hydrogen molality of approximately 56.55 $\cdot$(kg H2O)-1. Since this value is close to our upper limit of 42 $\cdot$(kg H2O)-1, we choose to adopt 56.55 $\cdot$(kg H2O)-1 as an upper limit instead.
@@ Appendix D Dilute Solution Assumptions in the DEW model
Our treatment of supercritical water as a dilute solution involves several assumptions that are implicit in our choice of standard states. A standard state is a set of conditions for a species with a fixed chemical composition, serving as a reference state for calculating thermodynamic properties. Often, the standard state is a hypothetical state that may not be physically possible but remains useful in extrapolating thermodynamic properties. We can relate the activities, molalities, and mole fractions through the definitions of standard states from the following equations, adopted from Sverjensky (2019).
@@
where ${\gamma}_{j}$ and ${m}_{j}$ refer to the activity coefficient and molality of species $j$ and ${\gamma}^{0}_{j}$ and ${m}^{0}_{j}$ refer to the standard state values (Anderson, 2005). We adopt a Henryan standard state (i.e., following Henry’s Law) known as the “ideal one molal state” for aqueous species. This is a commonly used hypothetical standard state that describes an ideal solution at infinite dilution with ${m}^{0}_{j}$ = 1.0. By definition, real aqueous species cannot physically exist in this state. Under this condition, since ${m}^{0}_{j}$ = 1.0 and ${\gamma}^{0}_{j}$ = 1.0 (see Figure 5 of Sverjensky (2019)), Equation D1 becomes,
@@
The species of interest in this paper (CH4, CO2, and CO) are neutral, nonpolar molecules, and while their activity coefficients may depart from unity, the ratio of their activity coefficients tends to cancel out (e.g., Barrett et al. 1988). As a result, the ratios of their activities can be approximated as,
$\frac{{a}_{j}}{{a}_{i}}\approx\frac{{m}_{j}}{{m}_{i}}$ (D4)
@@
where ${\lambda}_{\text{H}_{2}\text{O}}$ represents a rational activity coefficient. In the present system, water behaves ideally and ${\lambda}_{\text{H}_{2}\text{O}}$rightarrow 1.0 as ${X}_{\text{H}_{2}\text{O}}$rightarrow 1.0, and therefore,
@@
The results of Equations D4 and  D6 are hence substituted into Equations 2 and 4.
@@ Appendix E Generating the Pressure-Temperature Profiles
Care was taken in proximity to the liquid-vapor phase boundary, where a discontinuous jump in the adiabatic temperature gradient can produce errors in the data interpolation. Integration was stopped when the temperature was within 10 K of the phase boundary, and the profiles were then linearly interpolated (in logarithmic pressure space) towards the phase boundary. Warmer adiabats that do not intersect this boundary will have hydrogen in much lower concentrations in the upper atmosphere since in this case there is no cold trapping of water and the hydrogen will be fully miscible with supercritical water (Soubiran & Militzer, 2015), which is not consistent with the non-detection of water in K2-18 b’s atmosphere (Madhusudhan et al., 2023b).
@@ Appendix F Parameters for Equilibrium Calculations
The equilibrium constants and Henry’s Law constants in Equations 2, 4, 6, 8, and 10 are obtained with the DEW model, which uses a series of revised Helgeson-Kirkham-Flowers (HKF) equations of state for aqueous species (Shock & Helgeson, 1988, 1990; Sverjensky et al., 2014) and an equation of state by Zhang & Duan (2005) for water. For convenience, the computed values are then fitted with the logarithmic equation and associated values in Table F1. The equilibrium constants in Equations 2 and 4 are fitted on a linear grid of temperatures between 647.15 K - 1273.15 K at every 1 K following the lower limit, resembling the ocean temperature, at each surface pressure (1000 bar, 5000 bar, 10,000 bar).
[…]
Table F1: Summary of Key Chemical Reactions and Parameters to Calculate Equilibrium Constants
@@ A B C
[…]
Note. — This table presents the parameters for various chemical reactions. The constants A, B, and C are used in the logarithmic equation $\log(K)=A+\frac{B}{T}+C\log(T)$, where T refers to the absolute temperature.
Figure F.1: Logarithmicly fitted Henry’s Law constants for CH4, CO2, and CO across temperatures between 444.15 K to 623.15 K represented by solid lines, with extrapolated values between 624.15 K to 647.15 K represented by dashed lines. Pressures correspond to water-saturated conditions as shown in Figure 2a.
@@ G.1 P-T Profiles of Hydrogen-Water Mixtures
To address the challenges of modeling K2-18 b’s interior, we employed a geochemical model to investigate the chemistry of a hypothesized supercritical water ocean on K2-18 b. However, the existing DEW model is limited to water-rich fluids (see the assumptions in Appendix D) and cannot fully account for scenarios with higher hydrogen molalities. It is possible that K2-18 b possesses significantly higher hydrogen molalities than those modeled here. An upper limit on $m_{\text{H}_{2}}$ can be inferred from JWST-derived atmospheric abundances reported by Madhusudhan et al. (2023b). The volume mixing ratios of CH4 and CO2, each at approximately 1%, provide a lower bound on the bulk volatile content, thus constraining the upper limit on the hydrogen content. Using water as a proxy for all volatiles, this corresponds to an H2 molality upper limit of 1300 $\cdot$(kg H2O)-1. Such a high H2 molality would represent an extremely water-poor fluid, making it incompatible with the present model. To investigate these possible H2-rich scenarios, conventional gas-phase models (Yang & Hu, 2024) and/or molecular dynamics (MD) simulations such as those used in Gupta et al. (2024) and Cano Amoros et al. (2024) may be more appropriate.
@@
This potential limitation also affects our P-T profiles and the saturation boundary in Figure 2a, and consequently their intersections, where we performed Henry’s Law calculations. In this work, both the P-T profiles and the saturation boundary were constructed for pure water, or $X_{\text{H}_{2}\text{O}}$ = 100%. However, K2-18 b’s lower atmosphere has $X_{\text{H}_{2}\text{O}}$ $<$ 100%. In addition, it should be noted that we reference the critical point of pure water rather than that of an H2-H2O mixture. To estimate how different H2-H2O mixtures might influence the saturation boundary – and, by extension, the crossing temperatures and pressures – we analyzed the saturation boundary corresponding to each mole fraction of water used in this work. We used the Reference Fluid Thermodynamic and Transport Properties (REFPROP) program (Lemmon et al., 2018) to compare the difference in saturation temperatures up to the critical temperature of pure water. The results are summarized in Table G1. We find that these temperature changes are not significant enough to greatly impact our results, given other uncertainties such as the fluid oxidation state (see Appendix D). We note that this calculation does not include non-ideal gas behavior which becomes more prominent at higher pressures and consequently shifts the crossing temperature, however, changes in crossing temperature also do not matter much in the C-O-H system because solubility corrections applied to exsolved gases are similar (Figure F.1). Nevertheless, future work should consider the impact of non-ideal phases and incorporate P-T profiles and saturation boundaries that account for the effects of species other than water to achieve a more self-consistent model.
G.2 Carbon Dioxide Mixing Ratio
@@
We assume that gaseous calculations performed for the liquid-vapor interface (cloud bottom) can be directly compared to the JWST observations. However, one notable exception may be CO2, which can remain in a kinetically controlled balance with quenched CO and H2O at altitudes above the CO-CH4-H2O quench point. In that situation, the CO2 mixing ratio can often increase with altitude above the quench point, resulting in a CO2 mole fraction at higher altitudes in excess of expectations based on its deeper equilibrium abundance (Yu et al., 2021; Hu, 2021). Further disequilibrium chemistry gas-phase modeling may be needed to test this scenario.
G.3 Hydrogen-Water Immiscibility?
The pressures and temperatures we explored in this work partially overlap with a region of possible H2-H2O immiscibility between 5 kbar and 10 kbar, as shown in Figure 3 of Gupta et al. (2024). Instead of a homogeneous layer, H2-H2O immiscibility would lead to the formation of two phase-separated layers: an H2-rich layer on top of an H2O-rich layer. We performed additional calculations investigating how such phase separation could change our ocean temperature lower limit, which is currently given by the CH4/CO2 ratio. In this case, the equilibrium CH4/CO2 ratio is determined by the H2 molality and mole fraction of water in the H2O-rich layer on the bottom. If H2 is removed, this layer is expected to contain less dissolved hydrogen than a miscible H2/H2O layer, leading to a higher water mole fraction. Thus, we performed sensitivity tests self-consistently varying the mole fraction of water in our model between our highest water mole fraction, $X_{\text{H}_{2}\text{O}}$ = 99.6%, up to but not including $X_{\text{H}_{2}\text{O}}$ = 100.0% for our CH4/CO2 ratios at 5 kbar, the lowest pressure for possible immiscibility within our temperature range (Gupta et al., 2024). We found that our modeled CH4/CO2 ratios still agree with the atmospheric observations, given an ocean temperature lower limit of 710 K, when the mole fraction of water is less than $X_{\text{H}_{2}\text{O}}$ = 99.94% ($X_{\text{H}_{2}}$approx 0.06% or $m_{\text{H}_{2}}$approx 0.03 $\cdot$(kg H2O)-1). However, the lower temperature limit is likely to be immune to the effects of phase splitting, as it is determined by the most water-rich composition (Figure 3a). Miscibility can be expected at an H2 molality of 0.14 $\cdot$(kg H2O)-1. To decrease our current estimated lower ocean temperature limit below $\sim$ 710 K, the H2 molality would need to decrease by a factor of $\sim$5 in a phase-separated H2O-rich layer to achieve $X_{\text{H}_{2}\text{O}}$ = 99.94%. Meanwhile, to investigate how our upper limit on temperature may change due to higher water mole fractions in the phase-separated H2O-rich layer, we performed similar sensitivity tests for our CO/CO2 ratios at 10 kbar, the highest pressure of interest. Our CO/CO2 ratios are minimally changed since, as the water mole fraction approaches unity, the proportional increase in CO/CO2 reaches diminishing returns (Equation 4).
G.4 Supercritical-Liquid Water Adiabats
We find that a portion of our P-T profiles go directly from the supercritical phase to the liquid water phase, as seen by the leftmost adiabat for the 1 kbar pressure in Figure 2a. This implies a potential interior structure consisting of a hot liquid water layer above a supercritical water layer. However, it is important to note that our P-T profiles do not fully include climate considerations. As mentioned in Section 1, past climate models provide a strong case that surface water on K2-18 b would exist in the supercritical state. Innes et al. (2023) considered sub-Neptunes with cloud-free, pure H/He atmospheres receiving solar flux from M stars, and found that an atmospheric pressure of 10 bars will lead to a runaway greenhouse effect due to collision-induced absorption. As a result of the runaway greenhouse effect, any liquid water ocean would become supercritical. It has been suggested that surface water can remain liquid in the presence of clouds that provide the necessary albedo (Madhusudhan et al., 2023b). Recent studies argue that the high albedo required ($>$ 0.5) is both difficult to achieve for K2-18’s spectral type and inconsistent with observational data (Leconte et al., 2024), although such a high albedo cannot be entirely ruled out (see Cooke & Madhusudhan (2024)). These findings support previous interior models suggesting that water-rich sub-Neptunes with a runaway greenhouse effect will develop thick gaseous envelopes with supercritical water (Mousis et al., 2020). Therefore, when considering these climate models, we believe it is unlikely for surface water to remain in a liquid state. Our conclusion regarding habitability remains the same for either structure since all adiabats directly entering the liquid phase persist at temperatures above 500 K.
G.5 Species Abundances
Our approach allows us to discern the CH4/CO2 and CO/CO2 ratios that are consistent with observations. However, it is not yet applicable for calculating the mixing ratios of each species as this requires the detection of carbon-bearing molecules with a higher number of carbon atoms. If future observations detect molecules with more carbon atoms, we can begin to predict the absolute abundances with the DEW model.
Table G1: Comparison of H2-H2O Compositions and Critical Properties
[…]
(This work) (This work) (This work) (Seward & Franck, 1981) (Seward & Franck, 1981)
99.6% 0.2% 0.26 0.5% 647.45
90.0% 10.0% 8.79 9.0% 647.65
42.4%* 43.3%* 67.80 40.0% 656.45
Note. — *Derived from Yang & Hu (2024) and used in this work. We also note that not all of the H2O and H2 mole fractions used here sum to unity because of the addition of helium, while Seward & Franck (1981) conducted experiments with H2-H2O mixtures excluding helium. The maximum differences from saturated temperatures in this work are computed from the saturation boundary and critical temperature of pure water. Water clouds form at lower temperatures as the atmosphere becomes poorer in water vapor content